Skip to end of metadata
Go to start of metadata

TSE 02 was instrumented yesterday, some issues with the network but we may move an antenna soon.

The mote was not reporting but found the problem rather quickly. Measured for voltage - 0 volts. Turns out that the cable was chewed in two (gray PVC  cable, not the standard green serial cable)

replaced power supplies at V 06 and V 07 with 14 volt supplies

installed anti twist plate on 2m sonic at V 07 and got the DSM's on line again, but port 3 is not working (Licor). Possibly the fuse. Will look at later.

V06 has intermittent NAN's for the 20 m sonic.

installed USB stick at V 06 (was borrowed for another station back in Feb.)

Today (4/7) instrumented RSW 08, all data coming in, but voltage is low at DSM (10.9 volts), should be fine. Explored and found 2 towers which were erected, but not according to the spreadsheet. Will instrument those soon. 

2 Comments

  1. Remember that the Licor takes a 3A fuse.  We should have a bag of spares in the ops center.

    Feel free to update the spreadsheet when you find discrepancies.

     

  2. tse02 is on the net.  The problem was a typo for the PSK in the access point config.  Likewise the ISS ubiquiti station now links up, but I need to return to the site to plug the ubiquiti into the cradlepoint and replace the cellular Inernet wifi.